OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (5, 4) and (x, 7)

OpenStudy (anonymous):

Inverse variation can be modeled by the equation \[\Large y={k \over x}\] or \[\Large xy=k \] where k is some constant. If you have one point on the graph, you can plug in x and y and find k and then use it to find other values.

OpenStudy (anonymous):

In this case we want to find another point, but we only have one value, so if we find k, then we can go and find the missing x=value

OpenStudy (anonymous):

\[\frac{ 5 }{ 4} ?\]

OpenStudy (anonymous):

\[\Large 5={k \over 4}\] or \[\Large 5*4=k\] so what is the value of k?

OpenStudy (anonymous):

20 will be k ?

OpenStudy (anonymous):

Correct. So now that we know that, our second point is (x, 7) or in other words \[\Large 7={20 \over x}\] or \[\Large 7x=20\] So solve for x and you have your x-value

OpenStudy (anonymous):

2.8 ???

OpenStudy (anonymous):

Correct, it's a little weird for a question like this you usually just get an integer but that's correct.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

@doulikepiecauseidont

OpenStudy (anonymous):

Post this as a new thread and you'll be able to get even more feedback

OpenStudy (anonymous):

Oh wait nvm this is the same question

OpenStudy (anonymous):

From our problem solving steps we got \[\Large 7x=20\]

OpenStudy (anonymous):

lol is that correct ?

OpenStudy (anonymous):

So then we get \[\Large x={20 \over 7}\]

OpenStudy (anonymous):

And then 20= 14+6\[\Large \frac{14+6}{7}\]

OpenStudy (anonymous):

And you can separate the fractions to get \[\Large 2+\frac{6}{7}\]=\[\Large 2\frac{6}{7}\]

OpenStudy (anonymous):

And \[\Large \frac{14}{7}=2~ btw\]